On January 1, 2015, Pina Industries Inc. issued a $1,018,000, 10-year bond. The bond sold at 97, and paid 13% interest each January 1 and July 1. Pina called at 103 and cancelled the bond on January 1, 2020. Assume the company used the straight-line method of amortization.

1. Calculate the gain or loss on redemption.

2. Prepare the journal entry to record the early retirement of the bond.
Answer #1
Issue price of Bonds 987460 =1018000*0.97
Discount on issue 30540 =1018000-987460
Discount amortized for 5 years 15270 =30540*5/10
Unamortized discount 15270 =30540-15270
Carrying value of bonds on January 1, 2020 1002730 =1018000-15270
1
Bonds redemption price 1048540 =1018000*1.03
Less: Carrying value of bonds on January 1, 2020 1002730
Loss on redemption 45810
2
Debit Credit
Bonds payable 1018000
Loss on redemption of bonds 45810
       Discount on Bonds payable 15270
       Cash 1048540